# From Image Denoisers to Regularizing Imaging Inverse Problems: An Overview

### v1

@@ 5.3 Kurdyka–Łojasiewicz Property
Table 1: Summary of the properties of some convergent methods. By iterate convergence, we mean that the entire sequence converges to a point. For methods with residual convergence, they consider convergence of the form $\min_{l\leq k}\|x_{l+1}-x_{l}\|^{2}$. The KL property is used to transform the convergence of residuals to the convergence of iterates. When the denoiser takes the gradient-step form $D_{\sigma}=I-\nabla g_{\sigma}$, we denote by $L_{g}$ the Lipschitz constant of $g_{\sigma}$.
@@
In Table 1, we summarize the properties required for some recent provable PnP methods based on operator splitting convergence. Since the proof structure is quite similar for each of the methods given in Section 2, the constraints on the denoisers are also quite similar, and they even converge to critical points of the same functional. This is demonstrated in Figure 1, where for the deconvolution task with a fixed blur kernel and PnP denoiser, the reconstructions are all fairly similar. Figures 2 and 3 demonstrate the residual and peak signal-to-noise ratio (PSNR) convergence for a set of 10 images, again for the image deconvolution task. We observe that, as the theory suggests, the provable PnP methods exhibit decaying residuals and stable PSNR figures, whereas the non-provable DPIR method zhang2021plug does not demonstrate such convergence, while deteriorating in quality as the iterations continue.
@@
Figure 2: Residual convergence for deblurring on the CBSD10 dataset, with a uniform $9\times 9$ blur kernel and 3% additive Gaussian noise. We observe that while DPIR has slow residual convergence, the provable PnP methods all have a convergent behavior, often reaching their stopping criteria given by the change in objective value. In particular, the quasi-Newton PnP-LBFGS method converges very quickly within 100 iterations.
Figure 3: PSNR curves for deblurring on the CBSD10 dataset, with uniform $9\times 9$ blur kernel and 3% additive Gaussian noise. We observe that the non-provable DPIR method gradually decreases in PSNR at later iterations, eventually leading to instability. In contrast, the provable PnP methods all have stable convergence curves, reaching their stopping criteria.
@@ 6 Denoisers for Posterior Sampling
The use of denoisers as components in posterior sampling has gained considerable traction as the demand for uncertainty quantification in imaging grows. In many ill-posed inverse problems, the solution space is inherently ambiguous, and instead of recovering a single best estimate, it is often more informative to approximate the full posterior distribution $p(x|y)$. Denoisers provide a natural bridge for this, acting as powerful implicit priors that can be integrated into modern stochastic sampling schemes to draw samples from complex high-dimensional distributions.
@@ 6.1 The Bayesian Inversion Problem
When the forward model is linear with additive Gaussian noise, that is, $y=Kx+w$, where $w\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$, then the likelihood is given by $p(y|x)\propto\exp\left(-\frac{1}{2\sigma^{2}}\|y-Kx\|_{2}^{2}\right)$, providing a link with the fidelity term variational regularization. Sampling from this posterior is intractable if the prior is defined only implicitly through an image denoiser. Therefore, Monte Carlo Markov Chain (MCMC) algorithms are often used to generate samples from (an approximation to) the posterior by incorporating the image prior through an off-the-shelf pretrained image denoiser.
@@ 6.2 Plug-and-play Denoisers for Posterior Sampling
where $w$ is a standard Wiener process, $f(\cdot)$ defines the drift, and $g(\cdot)$ the diffusion strength. The time-reversed SDE is given by
@@
which guides the generative sampling process, where the time-dependent score $\nabla_{x}\log p_{t}(x)$ is approximated by a denoiser trained at varying noise scales song2021scorebased ; ho2020denoising . A particularly interesting special case is obtained by choosing $f(x,t)\equiv 0$ and a constant diffusion strength $g(t)$. In this case, the reverse SDE simplifies to
$\mathrm{d}x=\nabla_{x}\log p_{t}(x)\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}w,$
which is precisely the overdamped Langevin diffusion targeting the (instantaneous) data distribution with density proportional to $p_{t}(x)$. Equivalently, if the denoiser provides a score surrogate via Tweedie’s relation, one may write
$\mathrm{d}x\approx\Big{[}\tfrac{1}{\sigma^{2}}\big{(}D(x)-x\big{)}\Big{]}\mathrm{d}t+\sqrt{2}\,\mathrm{d}w,$
thereby exhibiting the plug-and-play interpretation in the unconditional setting. A first-order Euler–Maruyama discretization with step size $\delta>0$ yields the unadjusted Langevin algorithm (ULA), given by
$x_{k+1}=x_{k}+\delta\,\nabla_{x}\log p_{t}(x_{k})+\sqrt{2\delta}\,\boldsymbol{\epsilon}_{k},\qquad\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(0,I).$
@@
$x_{k+1}\approx x_{k}+\delta\,\tfrac{1}{\sigma^{2}}\big{(}D(x_{k})-x_{k}\big{)}+\sqrt{2\delta}\,\boldsymbol{\epsilon}_{k},$
the iteration refines samples by alternating deterministic drift along the (approximate) score with stochastic exploration pnp_ula_marcelo . This idea may be adapted to inverse problems kawar2022denoising , where the goal is to sample from a distribution consistent with both the learned image prior and the measurement model (see daras2024surveydiffusionmodelsinverse ; Chung2025 for recent surveys on the theory and applications of diffusion models for posterior sampling). A natural modification of the SDE to sample from the posterior distribution is by augmenting the drift with the likelihood score:
$\mathrm{d}x=\Big{[}s_{\theta}(x,t)+\nabla_{x}\log p(y|x)\Big{]}\mathrm{d}t+\sqrt{2}\,\mathrm{d}w,$
where $s_{\theta}(\cdot)$ is the learned score function corresponding to the image prior and $\nabla_{x}\log p(y|x)$ enforces data consistency. In the same special case $f(x,t)\equiv 0$, $g(t)\equiv\sqrt{2}$, this reduces to the overdamped Langevin diffusion for the measurement-informed target with density proportional to $p_{t}(x)\,p(y|x)$:
$\mathrm{d}x=\Big{[}\nabla_{x}\log p_{t}(x)+\nabla_{x}\log p(y|x)\Big{]}\mathrm{d}t+\sqrt{2}\,\mathrm{d}w.$
Under the linear Gaussian model $y=Kx+w$, $w\sim\mathcal{N}(0,\sigma_{w}^{2}I)$, the likelihood gradient takes the explicit form $\nabla_{x}\log p(y|x)=\tfrac{1}{\sigma_{w}^{2}}K^{*}(y-Kx)$, so that the conditional Langevin SDE reduces to
$\mathrm{d}x=\Big{[}\nabla_{x}\log p_{t}(x)+\tfrac{1}{\sigma_{w}^{2}}K^{*}(y-Kx)\Big{]}\mathrm{d}t+\sqrt{2}\,\mathrm{d}w.$
With a denoiser-based prior score surrogate, this can be approximated as $\mathrm{d}x\approx\Big{[}\tfrac{1}{\sigma^{2}}\big{(}D(x)-x\big{)}+\tfrac{1}{\sigma_{2}^{2}}K^{*}(y-Kx)\Big{]}\mathrm{d}t+\sqrt{2}\,\mathrm{d}w$. Discretizing again via Euler–Maruyama gives the Langevin-type update used in practice:
$x_{k+1}=x_{k}+\delta\Big{[}s_{\theta}(x_{k},t_{k})+\nabla_{x}\log p(y|x_{k})\Big{]}+\sqrt{2\delta}\,\boldsymbol{\epsilon}_{k},\quad\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(0,I).$
@@
$x_{k+1}\approx x_{k}+\delta\Big{[}\tfrac{1}{\sigma^{2}}\big{(}D(x_{k})-x_{k}\big{)}+\tfrac{1}{\sigma_{w}^{2}}K^{*}(y-Kx_{k})\Big{]}+\sqrt{2\delta}\,\boldsymbol{\epsilon}_{k},$
@@
Posterior sampling with denoisers has enabled uncertainty quantification in applications such as medical image reconstruction, compressive sensing, and computational microscopy. Multiple posterior samples allow practitioners to construct pixel-wise credible intervals and detect ambiguous regions that would otherwise be hidden by deterministic estimators. However, there are several practical challenges that remain to be addressed. Convergence guarantees for these denoiser-driven SDE samplers are still limited, especially when the denoiser is highly nonlinear and trained on finite data. The computational cost of generating many samples, which often requires thousands of iterative steps, can be computationally prohibitive. Active research seeks to develop more efficient discretizations, multi-scale models, or hybrid schemes that combine rapid MAP estimates with stochastic refinements to make these methods practical for large-scale problems. Nonetheless, the use of denoisers for posterior sampling illustrates the remarkable synergy between learned priors and stochastic inference for solving high-dimensional imaging inverse problems. By embedding denoisers within SDEs and MCMC updates, one can approximate complex posteriors that would be intractable otherwise, providing both high-fidelity reconstructions and rigorous uncertainty quantification. This paradigm represents an exciting frontier for solving ill-posed inverse problems in a principled, uncertainty-aware manner.
@@ 7 Conclusions and Outlook
We have surveyed the development of image denoising and the role that denoisers play in solving inverse problems through plug-and-play (PnP) methods. Beginning with classical denoising algorithms, we reviewed how modern learning-based denoisers can be seamlessly integrated into iterative schemes derived from variational regularization frameworks and proximal splitting algorithms. We discussed how PnP extends these algorithms by replacing proximal maps with powerful denoising operators, and explored related formulations such as Tweedie’s formula and the RED framework. Particular emphasis was placed on the mathematical conditions under which PnP methods converge, the constraints that must be imposed on denoisers to ensure stability, and practical considerations for deploying these techniques in real-world imaging settings.
@@
Looking forward, several promising research avenues emerge. First, a deeper theoretical understanding of PnP with non-expansive yet highly expressive denoisers could relax current restrictive assumptions while retaining convergence guarantees. Second, domain-adapted and multimodal denoisers have the potential to unlock PnP applications in emerging imaging modalities and dynamic acquisition settings. Third, the intersection of PnP with diffusion-based generative priors and self-supervised learning may yield reconstruction algorithms that are simultaneously more robust and data-efficient. Finally, scalable implementations capable of handling the high dimensionality and streaming nature of modern imaging data remain an important challenge, especially in time-critical domains such as medical imaging and remote sensing applications.

### v2

@@ 5.2 Backtracking for Lipschitz Control
Table 1: Summary of the properties of some convergent methods. By iterate convergence, we mean that the entire sequence converges to a point. For methods with residual convergence, they consider convergence of the form $\min_{l\leq k}\|x_{l+1}-x_{l}\|$. The KL property is used to transform the convergence of residuals to the convergence of iterates. Objective convergence denotes whether or not the cluster points are critical points of some computable function. The denoisers in the latter six methods use denoisers in gradient-step form $D_{\sigma}=\operatorname{Id}-\nabla g_{\sigma}$, and we denote by $L_{g}$ the Lipschitz constant of $g_{\sigma}$.
@@
In Table 1, we summarize the properties required for some recent provable PnP methods based on operator splitting convergence. To verify the convergence of the various provable PnP algorithms, we test them on a natural image deblurring task. We compare DPIR along with the latter five provable PnP algorithms in the table, where the denoiser is given by a gradient step denoiser $D_{\sigma}=\operatorname{Id}-\nabla g_{\sigma}$, where $g_{\sigma}(x)=\frac{1}{2}\|x-N_{\sigma}(x)\|^{2}$ is a pretrained DRUNet architecture zhang2021plug . To (approximately) satisfy the assumptions of the previous theorems, the Lipschitz constant of $\nabla g_{\sigma}$ is penalized to be less than 1 as in Section 5.1, and the activation functions are taken to be $\mathcal{C}^{2}$ and such that the neural network satisfies the KL property.
Since the proof structure is quite similar for each of the methods given in Section 2, the constraints on the denoisers are also quite similar, and they even converge to critical points of the same functional. This is demonstrated in Figure 1, where for the deconvolution task with a fixed blur kernel and PnP denoiser, the reconstructions are all fairly similar. Figures 2 and 3 demonstrate the residual and peak signal-to-noise ratio (PSNR) convergence for a set of 10 images, again for the image deconvolution task. We observe that, as the theory suggests, the provable PnP methods exhibit decaying residuals and stable PSNR figures, whereas the non-provable DPIR method zhang2021plug does not demonstrate such convergence, while deteriorating in quality as the iterations continue.
@@
Figure 2: Residual convergence for deblurring on the CBSD10 dataset, with a uniform $9\times 9$ blur kernel and 3% additive Gaussian noise. Each solid line represents one image. We observe that while DPIR has slow residual convergence, the provable PnP methods all have a convergent behavior, often reaching their stopping criteria given by the change in objective value. In particular, the quasi-Newton PnP-LBFGS method converges very quickly within 100 iterations.
Figure 3: PSNR curves for deblurring on the CBSD10 dataset, with uniform $9\times 9$ blur kernel and 3% additive Gaussian noise. Each solid line represents one image. We observe that the non-provable DPIR method gradually decreases in PSNR at later iterations, eventually leading to instability. In contrast, the provable PnP methods all have stable convergence curves, reaching their stopping criteria.
@@ 6 Denoisers for Posterior Sampling
The use of denoisers as components in posterior sampling has gained considerable traction as the demand for uncertainty quantification in imaging grows. In many ill-posed inverse problems, the solution space is inherently ambiguous. Instead of recovering a single best estimate, it is often more informative to approximate the full posterior distribution $p(x|y)$. Denoisers provide a natural bridge for this, acting as powerful implicit priors that can be integrated into modern stochastic sampling schemes to draw samples from complex high-dimensional distributions.
@@ 6.1 The Bayesian Inversion Problem
When the forward model is linear with additive Gaussian noise, that is, $y=Kx+w$, where $w\sim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I})$, then the likelihood is given by $p(y|x)\propto\exp\left(-\frac{1}{2\sigma_{w}^{2}}\|y-Kx\|_{2}^{2}\right)$, providing a link to the fidelity term within variational regularization. Sampling from this posterior is intractable if the prior is defined only implicitly through an image denoiser. Therefore, Monte Carlo Markov Chain (MCMC) algorithms are often used to generate samples from (an approximation to) the posterior by incorporating the image prior through an off-the-shelf pretrained image denoiser.
@@ 6.2 Plug-and-play Denoisers for Posterior Sampling
where $w$ is a standard Wiener process, $f(x,t)$ defines the drift, and $g(t)$ the diffusion strength. The time-reversed SDE is given by
@@
which guides the generative sampling process, where the time-dependent score $\nabla_{x}\log p_{t}(x)$ is approximated by a denoiser trained at varying noise scales song2021scorebased ; ho2020denoising . A particularly interesting special case is obtained by choosing $f(x,t)\equiv 0$ and a constant diffusion strength $g(t)\equiv 1$. In this case, the reverse SDE simplifies to
$\mathrm{d}x=-\nabla_{x}\log p_{t}(x)\,\mathrm{d}t+\mathrm{d}w,$ (46)
which is precisely the overdamped Langevin diffusion targeting the (instantaneous) data distribution with density proportional to $p_{t}(x)$. Equivalently, if the denoiser provides a score surrogate via Tweedie’s formula, one may write
$\mathrm{d}x\approx\left[-\tfrac{1}{\sigma_{t}^{2}}\left(D_{\sigma_{t}}(x_{t})-x_{t}\right)\right]\mathrm{d}t+\mathrm{d}w,$
thereby exhibiting the plug-and-play interpretation in the unconditional setting. Here, $D_{\sigma_{t}}$ is a denoiser trained or calibrated for noise level $\sigma_{t}$, and implicitly depends on time through the noise level $\sigma_{t}$, which plays the role of the diffusion variance at step $t$. A first-order Euler–Maruyama discretization with step size $\delta>0$ in backward time yields the unadjusted Langevin algorithm (ULA):
$x_{k+1}=x_{k}+\delta\,\nabla_{x}\log p_{t}(x_{k})+\sqrt{\delta}\,\boldsymbol{\epsilon}_{k},\qquad\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(0,I).$
@@
$x_{k+1}\approx x_{k}+\delta\,\tfrac{1}{\sigma_{k}^{2}}\big(D_{\sigma_{k}}(x_{k})-x_{k}\big)+\sqrt{\delta}\,\boldsymbol{\epsilon}_{k},$
the iteration refines samples by alternating deterministic drift along the (approximate) score with stochastic exploration pnp_ula_marcelo . This idea may be adapted to inverse problems kawar2022denoising , where the goal is to sample from a distribution consistent with both the learned image prior and the measurement model (see daras2024surveydiffusionmodelsinverse ; Chung2025 for recent surveys on the theory and applications of diffusion models for posterior sampling). A natural modification of the backward SDE 46 to sample from the posterior distribution is by augmenting the drift with the likelihood score:
$\mathrm{d}x=-\Big[\nabla_{x}\log p_{t}(x)+\nabla_{x}\log p(y|x)\Big]\mathrm{d}t+\mathrm{d}w.$
Under the linear Gaussian model $y=Kx+w$, $w\sim\mathcal{N}(0,\sigma_{w}^{2}I)$, the likelihood gradient takes the explicit form $\nabla_{x}\log p(y|x)=\tfrac{1}{\sigma_{w}^{2}}K^{\top}(y-Kx)$, so that the conditional Langevin SDE reduces to
$\mathrm{d}x=-\Big[\nabla_{x}\log p_{t}(x)+\tfrac{1}{\sigma_{w}^{2}}K^{\top}(y-Kx)\Big]\mathrm{d}t+\mathrm{d}w.$
With a denoiser-based prior score surrogate, this can be approximated as $\mathrm{d}x\approx-\Big[\tfrac{1}{\sigma_{t}^{2}}\big(D_{\sigma_{t}}(x)-x\big)+\tfrac{1}{\sigma_{w}^{2}}K^{\top}(y-Kx)\Big]\mathrm{d}t+\mathrm{d}w$. Discretizing again via Euler–Maruyama in reverse time gives the Langevin-type update used in practice:
$x_{k+1}=x_{k}+\delta\Big[s_{\theta}(x_{k},t_{k})+\nabla_{x}\log p(y|x_{k})\Big]+\sqrt{\delta}\,\boldsymbol{\epsilon}_{k},\quad\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(0,I).$
@@
$x_{k+1}\approx x_{k}+\delta\Big[\tfrac{1}{\sigma_{k}^{2}}\big(D_{\sigma_{k}}(x_{k})-x_{k}\big)+\tfrac{1}{\sigma_{w}^{2}}K^{\top}(y-Kx_{k})\Big]+\sqrt{\delta}\,\boldsymbol{\epsilon}_{k},$
@@
Posterior sampling with denoisers has enabled uncertainty quantification in applications such as medical image reconstruction, compressive sensing, and computational microscopy. Multiple posterior samples allow practitioners to construct pixel-wise credible intervals and detect ambiguous regions that would otherwise be hidden by deterministic estimators. However, there are several practical challenges that remain to be addressed. Convergence guarantees for these denoiser-driven SDE samplers are still limited, especially when the denoiser is highly nonlinear and trained on finite data. The computational cost of generating many samples, which often requires thousands of iterative steps, can be computationally prohibitive. Active research seeks to develop more efficient discretizations learning_disc_icml , latent diffusion models latent_diffusion_cvpr , or hybrid schemes that combine rapid MAP estimates with stochastic refinements NEURIPS2024_2f46ef57 to make these methods practical for large-scale problems. Nonetheless, the use of denoisers for posterior sampling illustrates the remarkable synergy between learned priors and stochastic inference for solving high-dimensional imaging inverse problems.
@@ 7 Conclusions and Outlook
In this chapter, we have surveyed the development of image denoising and the role that denoisers play in solving inverse problems through plug-and-play (PnP) methods. Beginning with classical denoising algorithms, we reviewed how modern learning-based denoisers can be seamlessly integrated into iterative schemes derived from variational regularization frameworks and proximal splitting algorithms. We discussed how PnP extends these algorithms by replacing proximal maps with powerful denoising operators, and explored related formulations such as Tweedie’s formula and the RED framework. Particular emphasis was placed on the mathematical conditions under which PnP methods converge, the constraints that must be imposed on denoisers to ensure stability, and practical considerations for deploying these techniques in real-world imaging settings.
@@
Looking forward, several promising research avenues emerge. First, a deeper theoretical understanding of PnP with non-expansive yet highly expressive denoisers could relax current restrictive assumptions while retaining convergence guarantees. Second, domain-adapted and multimodal denoisers have the potential to unlock PnP applications in emerging imaging modalities and dynamic acquisition settings. Third, the intersection of PnP with diffusion-based generative priors and self-supervised learning may yield reconstruction algorithms that are simultaneously more robust and data-efficient. By embedding denoisers within SDEs and MCMC updates, one can approximate complex posteriors that would be intractable otherwise, providing both high-fidelity reconstructions and rigorous uncertainty quantification. This paradigm represents an exciting frontier for solving ill-posed inverse problems in a principled, uncertainty-aware manner. Finally, scalable implementations capable of handling the high dimensionality and streaming nature of modern imaging data remain an important challenge, especially in time-critical domains such as medical imaging and remote sensing applications.